# On the numerical stability of sketched GMRES

### v1

@@ Abstract
We perform a backward stability analysis of preconditioned sketched GMRES [Nakatsukasa and Tropp, SIAM J. Matrix Anal. Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff. Under this condition, we show that sketched GMRES is backward stable as long as the condition number of $B_{1:i}$ is not too large. Under additional assumptions, we then show that the stability of a restarted implementation of sketched GMRES can be independent of the condition number of $B_{1:i}$, and restarted sketched GMRES is backward stable. We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically. We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
@@ 1 Introduction
In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [10] for solving nonsymmetric and nonsingular linear systems of the form
$Ax=b,\enspace A\in\mathbb{R}^{n\times n}$ (1)
@@
Although standard GMRES [12] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
@@
Sketched GMRES was proposed in [10] as an attractive alternative to GMRES that exploits the technique of randomized sketching [9] to avoid working with an orthonormal basis. Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
It has been observed in numerical experiments in [10] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [7] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly. Indeed, numerical experiments in [7] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [10, 7]. Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$. We then provide a backward stability result for a restarted version of sketched GMRES, which under certain assumptions, allows us to recover the numerical stability of sketched GMRES when $\kappa(Z_{1:i})$ grows large, suggesting we should monitor the conditioning of the basis throughout the iterations and restart when necessary. We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
[…]
@@
The remainder of this paper is organized as follows. In Section 2, we give an overview of GMRES and the MOD-GMRES framework [3] for performing a backward stability analysis of different GMRES variants. We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5. Numerical experiments are presented in Section 6.
@@ 2 Preconditioned GMRES
There are many algorithmic variants of GMRES arising from different implementation options, such as the choice of preconditioner (if any), or the choice of orthogonalization method used within the Arnoldi procedure. Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES. In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [3], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants. The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
Without loss of generality, we assume $x^{(0)}=0$ in this framework and our analysis of sGMRES in Section 4. By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis. For example, well-known results from [5] were derived in [3] using the framework. In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable. Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [11], was also recovered using the framework in [3].
@@ 3 Preconditioned Sketched GMRES
for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$. The sketching matrix $S$ needs to be drawn at random to ensure the relation (5) holds with a high probability. In practice, the sketched least squares problem (4) is computed by solving a triangular linear system
@@
It is clear from experimental observations made in [10, 7] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES. The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, where each new Krylov vector is orthogonalized against the previous $t$ vectors. Alternatively, the sketch-and-select method [7] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
@@ 4 Backward stability of preconditioned sketched GMRES
In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [3]. The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [8]. Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [3], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$. The definition of the key dimension is as follows:
$\sigma_{\min}(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq\omega_{\kappa(r,W)}\bigl{\|}\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (9)
@@
We first set up the least squares problem by constructing the matrix $\hat{C}_{1:k}$, and the vector $g:=SM_{L}^{-1}b$. We assume that
$\hat{C}_{1:k}$ $=S\hat{W}_{1:k}+\Delta C_{1:k},\quad\|\Delta C_{j}\rVert\leq\varepsilon_{S}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k,$ (11)
@@
and furthermore, that the computed solution $\hat{y}^{(k)}$ of the least squares problem in Step 2 satisfies
$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:k}+\Delta\hat{C}_{1:k})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq k+1,\end{split}$ (13)
with $\varepsilon_{sls}\leq O(\bm{u})$ from [8, Theorem 20.3]. Substituting $\hat{C}_{1:k}$ and $\hat{g}$ in (13) with (11) and (12), we have
$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\big{\|}S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k})y\big{\rVert},\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k.\end{split}$ (14)
We can now prove a bound on the norm of the sketched residual at iteration $k$, i.e., the norm of the vector
@@
$\sigma_{\min}(\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (16)
@@
We follow the proof in [3] and employ [11, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [11, Theorem 2.4] to the least squares problem (14), we have
@@
We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [3, (3.17)], which can simplify (19) as
@@
It remains to bound $\sigma_{\min}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}$ and $\phi$ in (21). By the assumption (16), and the estimation of $\|\Delta\tilde{g}\rVert$ and $\|\Delta\tilde{C}_{1:k}\rVert$ shown in (14), we have
@@
For $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and $\hat{y}^{(k)}\in\mathbb{R}^{k}$, we assume that
$\hat{x}^{(k)}=\hat{Z}_{1:k}\hat{y}^{(k)}+\Delta x^{(k)},\quad\|\Delta x^{(k)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:k}\rVert\|\hat{y}^{(k)}\rVert$ (27)
with $\varepsilon_{x}\leq O(\bm{u})$ from [8].
@@
The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
@@
Furthermore, from [10, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
@@
It remains to bound $\|r^{(k)}_{S}\rVert$ and $\|\hat{y}^{(k)}\rVert$. To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17). From [1, Corollary 2.2], the condition numbers of the sketched matrices satisfy
@@
$\sigma_{\min}(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\|S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\rVert$
@@
If the preconditioners satisfy $M_{L}=M_{R}=I$, then the relative backward error can be bounded as
@@
Let $\varepsilon_{*}$ denote all the $\varepsilon$ with any subscript mentioned in this section. Considering $\varepsilon_{*}\leq O(\bm{u})$, note that the assumption (28) in Theorem 1 can be replaced by
@@
from (7). Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [3, Theorem 3.1].
@@
In has been observed in numerical experiments in [6, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight. Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
5 Adaptively restarting sGMRES
Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process, and can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [3, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).

### v2

@@ Abstract
We perform a backward stability analysis of preconditioned sketched GMRES [Nakatsukasa and Tropp, SIAM J. Matrix Anal. Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff. Under this condition, we demonstrate that the stability of the sketched GMRES is mainly affected by the condition number of $B_{1:i}$, which can be mitigated by using the restarting technique. We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically. We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
@@ 1 Introduction
In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [NT2024] for solving nonsymmetric and nonsingular linear systems of the form
$Ax=b,\kern 5.0ptA\in\mathbb{R}^{n\times n}$ (1)
@@
Although standard GMRES [SaadSchultz1986] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
@@
Sketched GMRES was proposed in [NT2024] as an attractive alternative to GMRES that exploits the technique of randomized sketching [martinsson2020randomized] to avoid the high cost of constructing an orthonormal basis. Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
It has been observed in numerical experiments in [NT2024] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [guttel2024sketch] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly. Indeed, numerical experiments in [guttel2024sketch] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [NT2024, guttel2024sketch]. Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$. We then provide a backward stability result for a restarted version of sketched GMRES, which under certain assumptions, allows us to recover the numerical stability of sketched GMRES when $\kappa(Z_{1:i})$ grows large, suggesting we should monitor the conditioning of the basis throughout the iterations and restart when necessary. We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
We briefly outline the notation and some basic definitions used throughout this paper. We use MATLAB indexing to denote submatrices. For example, we use $X_{1:i}$ to denote the first $i$ columns of $X$. We use superscripts to denote iteration numbers. For example, $x^{(i)}$ denotes the approximate solution of $Ax=b$ at the $i$-th iteration of GMRES or sketched GMRES. We use $\|\cdot\rVert$ to denote the vector $2$ norm and $\bigl\|\cdot\bigr\rVert_{\mathsf{F}}$ to denote the Frobenius norm. We use $\bm{u}$ to denote the unit roundoff. We also use $\hat{\cdot}$ to denote computed quantities. For example, $\hat{x}$ denotes the computed value of the vector $x$ in finite precision. The (normwise) backward error of $\hat{x}$ is defined as
@@
The remainder of this paper is organized as follows. In Section 2, we give an overview of GMRES and the MOD-GMRES framework [BHMV2024] for performing a backward stability analysis of different GMRES variants. We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5. Numerical experiments are presented in Section 6.
@@ 2 Preconditioned GMRES
There are many algorithmic variants of GMRES arising from different implementation options, such as the choice of preconditioner (if any), or the choice of orthogonalization method used within the Arnoldi procedure. Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES. In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [BHMV2024], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants. The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
Without loss of generality, we assume $x^{(0)}=0$ in this framework and our analysis of sGMRES in Section 4. By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis. For example, well-known results from [drkovsova1995numerical] were derived in [BHMV2024] using the framework. In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable. Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [paige2006modified], was also recovered using the framework in [BHMV2024].
@@ 3 Preconditioned Sketched GMRES
for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$. In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (5) with high probability. Additionally, $S$ must be constructed without prior knowledge of the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$ and should typically be a valid embedding for a broad subspace, specifically $\mathbb{R}^{n\times i}$. In practice, the sketched least squares problem (4) is computed by solving a triangular linear system
@@
It is clear from experimental observations made in [NT2024, guttel2024sketch] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES. The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, as summarized in Algorithm 3, where each new Krylov vector is orthogonalized against the previous $t$ vectors. Alternatively, the sketch-and-select method [guttel2024sketch] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
0: A matrix $A\in\mathbb{R}^{n\times n}$, a truncated parameter $t$, the basis $B_{1:i-1}$ generated by the first i1i-1 truncated Arnoldi iterations.
0: The basis $B_{1:i}$.
[…]
2:for $l=\max(1,i-t+1):i$ do
3:  $B_{i}=B_{i}-B_{l}(B_{l}^{\top}B_{i})$.
4:end for
[…]
Algorithm 3 The $i$-th iteration of the truncated Arnoldi process
@@ 4 Backward stability of preconditioned sketched GMRES
In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [BHMV2024]. The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [H2002]. Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [BHMV2024], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$. The definition of the key dimension is as follows:
[…]
@@
We first set up the least squares problem by constructing the matrix $\hat{C}_{1:i}$, and the vector $g:=SM_{L}^{-1}b$. We assume that
$\hat{C}_{1:i}$ $=S\hat{W}_{1:i}+\Delta C_{1:i},\quad\|\Delta C_{j}\rVert\leq\varepsilon_{S}\|S\hat{W}_{j}\rVert,\quad\forall j\leq i,$ (11)
@@
and furthermore, that the computed solution $\hat{y}^{(i)}$ of the least squares problem in Step 2 satisfies
$\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:i}+\Delta\hat{C}_{1:i})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq i+1,\end{split}$ (13)
with $\varepsilon_{sls}\leq O(\bm{u})$ from [H2002, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
$\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\big\|S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:i}+\Delta\tilde{C}_{1:i})y\big\rVert,\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{W}_{j}\rVert,\quad\forall j\leq i.\end{split}$ (14)
We can now prove a bound on the norm of the sketched residual at the key dimension $k$, i.e., the norm of the vector
@@
[…]
@@
We follow the proof in [BHMV2024] and employ [paige2006modified, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [paige2006modified, Theorem 2.4] to the least squares problem (14), we have
@@
We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [BHMV2024, (3.17)], which can simplify (19) as
@@
[…]
@@
For $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and $\hat{y}^{(i)}\in\mathbb{R}^{i}$, we assume that
$\hat{x}^{(i)}=\hat{Z}_{1:i}\hat{y}^{(i)}+\Delta x^{(i)},\quad\|\Delta x^{(i)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:i}\rVert\|\hat{y}^{(i)}\rVert$ (27)
with $\varepsilon_{x}\leq O(\bm{u})$ from [H2002].
@@
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied. Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$. If it further holds that
@@
Note that $\hat{b}\in\mathbb{R}^{n}$, and we can establish the existence of $k$, since the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is satisfied when $k=n$.
Then we only need to prove (29). The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
@@
Furthermore, from [NT2024, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
@@
It remains to bound $\|r^{(k)}_{S}\rVert$ and $\|\hat{y}^{(k)}\rVert$. To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17). From [BG2022, Corollary 2.2], the condition numbers of the sketched matrices satisfy
@@
[…]
@@
Theorem 1 indicates that (29) and (30) are satisfied with high probability. This is due to the $\epsilon$-subspace embedding, which generally meets the condition (5) with high probability. Practically, the sketching method rarely fails, and slightly increasing the embedding dimension can further improve its reliability. In the worst case, $\epsilon$ may approach $1$ and the constant $(1+\epsilon)/(1-\epsilon)$ may be relatively large. For this reason, it is distracting to quantify the failure probabilities; see details in [MT2020].
Remark 2.
Let $\varepsilon_{*}$ denote all the $\varepsilon$ with any subscript mentioned in this section. Assuming $\varepsilon_{*}\leq O(\bm{u})$, we simplify Theorem 1 for the case without preconditioning, i.e., $M_{L}=M_{R}=I$ and $\kappa(\hat{Z}_{1:k})=\kappa(\hat{B}_{1:k})$: if it holds for the key dimension $k$ that
$O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,\qquad O(\bm{u})\kappa(\hat{B}_{1:k})\leq 1/2,$ (39)
then the relative backward error can be bounded as
@@
Considering $\varepsilon_{*}\leq O(\bm{u})$, note that the assumption (28) in Theorem 1 can be replaced by
@@
from (7). Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [BHMV2024, Theorem 3.1].
@@
It has been observed in numerical experiments in [guttel2024sketch, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight. Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
5 Adaptive restarting sGMRES
Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process as follows:
  1. 1.
    Compute the residual $r^{(i)}=b-Ax^{(i)}$.
  2. 2.
    Solve $Ad^{(i)}=r^{(i)}$, which is computed by Algorithm 1.
  3. 3.
    Update the solution $x^{(i+1)}=x^{(i)}+d^{(i)}$.
Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [BHMV2024, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).